# Cosmological constant as an integration constant

### v3

@@ Abstract
The discrepancy between the observed value of the cosmological constant (CC) and its expected value from quantum field theoretical considerations motivates the search for a theory in which the CC is decoupled from the vacuum energy. In this article, we consider theories in which the CC is regarded as an integration constant. These theories include trace-free Einstein gravity, theories constructed from the Codazzi equation (which includes Cotton gravity and a gauge-gravity inspired theory), and conformal Killing gravity. We remark on a recent debate regarding the viability of Cotton gravity and find that while the Codazzi equation is indeed underdetermined, the solutions of the Codazzi equation trivialize to $\lambda g_{ab}$ on generic backgrounds, and that in principle, one can close the system with the divergence-free condition and an appropriate choice of initial data. We also propose a full variational principle (full in the sense that variations in all variables are considered) for each of the aforementioned theories that can incorporate the matter sector; in this manner, we can obtain the trace-free Einstein equations without a unimodular constraint.
@@
Introduction — Since Einstein introduced the cosmological constant (CC) to his equations in 1917 [1], the debate about its nature has never stopped. Before the discovery of the accelerating expansion of the universe in 1999 [2, 3], the main consideration was why the CC is not exactly zero (See, for example, the famous review by Weinberg[4]). In the aftermath of the discovery of the accelerating expansion of the universe, the problem became more acute. The notion of dark energy introduced to account for this effect behaves like the CC. Indeed, in most analyses in cosmology, such as that of the cosmic microwave background, the CC is commonly invoked in the modeling. If true, then the observed value for the CC is tiny compared with the microscopic interpretation of the CC a la quantum field theory. Why is the CC not only nonzero but also so tiny? Such a seeming dilemma further motivates the reconstruction of Einstein gravity.
@@
Such a reformulation should somehow decouple the CC from microscopic physics. One strategy is to seek a formulation of Einstein gravity in which the CC arises as a constant of integration. At the level of the field equations, there are several proposals in the literature. A proposal originating from Einstein [5] is that of the trace-free Einstein equations, which are obtained in unimodular gravity [6, *vanderBij:1981ym, *Buchmuller:1988wx, *Buchmuller:1988yn, *Henneaux:1989zc, 4] (see [11, *Alvarez:2023utn] for recent overviews of the topic), but can be considered on their own without the constraint on the metric determinant [13, 14]. Another proposal by Cook and Chen [15, 16], inspired by a set of third-order equations developed by Kilmister, Newman [17, *Kilmister:1961], and later by Yang [19] is to construct a gravity theory from the Bianchi identities and derivatives of the energy-momentum tensor. We focus primarily on the aforementioned approaches, though we mention here other proposals in which one may obtain the unconstrained Einstein field equations with the CC arising as an integration constant, such as the thermodynamical approach of Padmanabhan [20, 21] (see also [22]), a modified metric-affine theory [23, *Tapia:1998qt, *Tapia:1998as], and contributions from a three-form action  [^1] However, we note that in this case the vacuum energy still contributes to the cosmological term. [26, 27]. More recently, Harada proposed two theories of gravity, termed Cotton gravity [28] and the so-called conformal Killing gravity [29, 30]. Cotton gravity is constructed from the Cotton tensor, and it has been shown that one can rewrite the field equations of Cotton gravity as a Codazzi equation $\nabla_{[a}C_{b]c}=0$ [31, 32]; in this regard, Cotton gravity is in the same class as the earlier theory of Cook and Chen, as we will show that the field equations for the latter can also be reformulated as a Codazzi equation. Conformal Killing gravity, on the other hand, has a different structure altogether, as the field equations have the form of the defining equation for a gradient conformal Killing tensor.
@@
Cotton gravity was recently the focus of a debate [33, 34, 35, 36] over its viability as a physical theory, as it was discovered that the symmetry-reduced equations of Cotton gravity are underdetermined. It is been argued in [34, 36] that while the equations of Cotton gravity are indeed underdetermined in cases of high symmetry (for instance, cosmological or spherically symmetric spacetimes), this is not necessarily the case in generic spacetimes. Similar concerns apply to the theory of Cook and Chen, as they share the same Codazzi structure. Another concern that applies to Codazzi type theories and conformal Killing gravity, is the fact that they are only defined at the level of the field equations; these theories suffer from the lack of a complete variational principle, as the variational principles proposed so far require holding some of the variables (such as the spacetime metric $g_{ab}$) fixed. One difficulty is that the energy-momentum tensor is obtained from variations of the metric tensor in the matter sector, but derivatives of the energy-momentum tensor appear in the field equations. The primary aim of this article is to shed light on the debate over the Codazzi equation and to propose a variational principle (including matter) for the trace-free Einstein equations, the Codazzi equation, and the conformal Killing equation without introducing additional constraints.
@@
Trace-free Einstein equations — The trace-free Einstein equations may be obtained by assuming that the determinant of the metric is held fixed under the variation of the action; in this form, the theory is called unimodular gravity. However, the equations themselves can be written without assuming such a constraint [13, 14]:
@@
Bianchi gravity — The works [15, 16] consider a field theory in which the connection and curvature tensors are treated as analogues of the respective gauge field and field strength tensor of electrodynamics, yielding:
@@
Equation (9) has the form of a Codazzi equation [31, 32], which is the subject of recent interest in the Cotton gravity program [28].
Cotton gravity — Cotton gravity is based on the Cotton tensor:
@@
one can write the governing equation for Cotton gravity as:
@@
“Conformal” Killing gravity — The equations for the so-called conformal Killing gravity theory proposed in [29] may be written in the form [30]:
@@
where $u_{a}$ is a one-form; $K_{ab}$ is a symmetric Killing tensor when $u_{a}=0$. However, as pointed out in [37] a simple redefinition $\bar{K}_{ab}:=6K_{ab}-g_{ab}K$ reveals that any nontrivial solution of Eq. (16) requires the existence of a Killing tensor given by $\bar{K}_{ab}$. It follows that the solutions of Eq. (16) either trivialize to $K_{ab}=-\lambda g_{ab}$, or admit a Killing tensor. For this reason, it is perhaps more appropriate to refer to this theory as “Killing tensor gravity”, but we retain here the common terminology in the literature. Moreover, we note that perturbations of Killing tensor solutions satisfying $K_{ab}\neq-\lambda g_{ab}$ must also admit a Killing tensor; since our universe (being inhomogeneous and anisotropic on small scales) does not have the required symmetries to admit a global Killing tensor, we regard such solutions as unphysical.
The Codazzi equation — Equations (9) and (14) have the form of a Codazzi equation:
@@
where $C_{ab}=C_{ba}$ is termed a Codazzi tensor, following the terminology of [31, 32]. One can perhaps imagine a class of theories defined by an equation of the form in Eq. (18). Now there is some debate over whether equations of this type are predictive [33, 34, 35, 36]; in particular, it has been claimed in [35] that the Codazzi equations are underdetermined, and in a responding comment it was claimed that this only occurs for highly symmetric spacetimes [36]. In the following, we attempt to shed some light on this dispute.
@@
Of course, Eq. (20) is underdetermined due to the term $\nabla_{a}{\nabla}_{c}C{}^{cb}$. In an orthonormal basis, the derivative $\partial^{2}_{t}C{{}_{t}}^{b}$ disappears from Eq. (20), so that the time evolution of the components $C{{}_{t}}^{b}$ is ill-defined—the equations are indeed underdetermined. However, Eq. (20) reveals a straightforward resolution; one should supply the Codazzi equation with a constraint equation, and a natural one is the divergence-free condition $\nabla_{a}C^{ab}=0$. Of course, an attempt to directly solve the Codazzi equations supplemented with the $\nabla_{a}C^{ab}=0$ may still lead to ambiguities, but such ambiguities correspond to freedom in choosing initial data for $C_{ab}$ and $\partial_{t}C_{ab}$ for the equivalent system:
@@
Now Eq. (20) follows from Eq. (18). It is perhaps appropriate to consider the application of the D’Alembertian to the following quantity:
@@
Note that the right-hand side contains at most first derivatives of $C_{ab}$ (and also that the divergence-free condition removes the first term). One can then regard Eq. (24) as a wave equation for $Q_{cde}$, with $C_{ab}$ being an independent variable satisfying Eq. (20) and the constraint (24). Now consider initial data consistent with Eq. (18), so that $Q_{cde}=0$ (but not its derivatives). For nonvanishing curvature, the right-hand side will generically produce a source unless $C_{ab}=\lambda g_{ab}$ on a Cauchy hypersurface and $\nabla_{c}C_{ab}=0$. This indicates that if $C_{ab}\neq\lambda g_{ab}$ and $\nabla_{c}C_{ab}\neq 0$ satisfies Eq. (24) on a Cauchy surface, the time evolution of $C_{ab}$ under Eq. (20) will in general generate violations of Eq. (24). There may also be special cases (highly symmetric spacetimes, for instance) in which Eq. (18) may admit nontrivial solutions $C_{ab}\neq\lambda g_{ab}$, but one cannot expect the existence of solutions to Eq. (18) for generic initial data. It should be noted that Eq. (24) and the conclusions outlined in the previous paragraph are independent of the constraint $\nabla_{a}C^{ab}=0$.
In short, Eq. (20) indicates that the quantity $Q_{ab}{{}^{c}}$ evolves according to Eq. (24), and that generically, the condition $Q_{ab}{{}^{c}}=0$ can only be maintained for initial data also satisfying $C_{ab}=\lambda g_{ab}$ and $\nabla_{c}C_{ab}=0$, where $\lambda$ is an arbitrary constant; other solutions will likely require background geometries with special properties, as argued in [34, 36]. We find that both sets of authors, [33, 35] and [34, 36], present valid points: theories based on the Codazzi tensors are indeed underdetermined, but the space of solutions is nonetheless highly restricted in generic settings.
Variational principle — The variational principles presented in the literature for the trace-free Einstein equations [4] and theories admitting a Codazzi equation formulation (Bianchi gravity [15, 16] and Cotton gravity [28]) typically require holding certain quantities fixed; for the trace-free Einstein equations, the metric determinant is held fixed (corresponding to an additional constraint), and for the Codazzi equation, the metric is held fixed (as variations in the metric will generally introduce additional unwanted constraints). We therefore seek a general variational principle that yields the desired field equations (including contributions from the matter sector) without unwanted constraints under a full variation—that is, a variation with respect to all variables appearing in the action.
@@
where $\underline{R}_{ab}:=R_{ab}-\tfrac{1}{4}Rg_{ab}$ and $\underline{Z}_{ab}:=R_{ab}-\tfrac{1}{4}Zg_{ab}$. We emphasize that no constraint on the metric determinant is required.
@@
where $\phi$ is a matter field (assumed to be nonspinorial and minimally coupled), $\chi$ is an extra field that is equal to $\phi$ on shell, $h^{ab}$ and $J$ are auxiliary fields, $\zeta$ is a Lagrange multiplier [^3] In particular, one assumes $\zeta$ to be defined so that $J\cdot\zeta\cdot J$ is a scalar., and the following have been defined:

### v4

@@ Abstract
The discrepancy between the observed value of the cosmological constant (CC) and its expected value from quantum field theoretical considerations motivates the search for a theory in which the CC is decoupled from the vacuum energy. In this article, we consider the viability of theories in which the Einstein equations are recovered (without additional constraints) and in which the CC is regarded as an integration constant. These theories include trace-free Einstein gravity, theories constructed from the Codazzi equation (which includes Cotton gravity and a gauge-gravity inspired theory), and conformal Killing gravity. We remark on a recent debate regarding Cotton gravity and find that while the Codazzi equation of that theory is indeed underdetermined, the solutions of the Codazzi equation trivialize to $\lambda g_{ab}$ on generic backgrounds, and that in principle, one can close the system with the divergence-free condition and an appropriate choice of initial data. We also propose a full variational principle (full in the sense that variations in all variables are considered) for each of the aforementioned theories that can incorporate the matter sector; in this manner, we can obtain the trace-free Einstein equations without a unimodular constraint. The resulting actions require additional (auxiliary) fields and are therefore only expected to be effective, but they may provide a useful starting point in bottom up approaches to constructing more fundamental theories.
keywords:
Cosmological constant problem , Cotton gravity , Codazzi tensor , Trace-Free Einstein equations , Conformal Killing gravity
journal: Physics Letters B
[…]
[aff1]organization=Leung Center for Cosmology and Particle Astrophysics, National Taiwan University, addressline=No.1 Sec.4, Roosevelt Rd., city=Taipei, postcode=10617, state=, country=Taiwan, R.O.C.
1 Introduction
@@
Since Einstein introduced the cosmological constant (CC) to his equations in 1917 [22], the debate about its nature has never stopped. Before the discovery of the accelerating expansion of the universe in 1999 [47; 46], the main consideration was whether the CC is exactly zero (See, for example, the famous review by Weinberg [58]). In the aftermath of the discovery of the accelerating expansion of the universe, the problem became more acute. The notion of dark energy introduced to account for this effect behaves like the CC. Indeed, in most analyses in cosmology, such as that of the cosmic microwave background, the CC is commonly invoked in the modeling. If true, then the observed value for the CC, fixed by observations of the accelerating expansion of the universe [14], is many orders (up to 120) of magnitude smaller than contributions from vacuum fluctuations arising from quantum field theoretic considerations. The discrepancy between the vacuum energy and the observed value for the CC requires an extreme degree of fine tuning. On the other hand, the Einstein field equations have been well-established on astrophysical scales, even in the strong gravity regime. This state of affairs strongly motivates a modification or reconstruction of Einstein gravity.
In this article, we consider one strategy, discussed in Section VII of [58], in which one seeks a formulation of Einstein gravity in which the CC arises as a constant of integration in a manner independent of vacuum fluctuations, avoiding the need for fine tuning in this regard. Though the most well-known of these approaches is that of unimodular gravity [4; 9; 10; 11; 29; 58] (see [12; 1] for recent overviews of the topic), which can be formulated at the level of the action, unimodular gravity introduces a constraint on the metric determinant that breaks general covariance. On the other hand, at the level of the field equations, there are several proposals in the literature that do not require such a constraint; our main aim in this article will be to evaluate the viability of these approaches.
@@
A proposal originating from Einstein [23] is that of the trace-free Einstein equations, which are recovered in unimodular gravity, but can be considered on their own without the unimodular constraint [24; 13]. Another proposal by Cook and Chen [17; 14], inspired by a set of third-order equations developed by Kilmister, Newman [31; 32], and later by Yang [59] is to construct a gravity theory from the Bianchi identities and derivatives of the energy-momentum tensor. We focus primarily on the aforementioned approaches, though we mention here other proposals in which one may obtain the unconstrained Einstein field equations with the CC arising as an integration constant, such as the thermodynamical approach of Padmanabhan [42; 41] (see also [18]), a modified metric-affine theory [55; 54; 56], and contributions from a three-form action  [^1] However, we note that in this case the vacuum energy still contributes to the cosmological term. [21; 28].
More recently, Harada proposed two theories of gravity, termed Cotton gravity [26] and the so-called conformal Killing gravity [27; 35]. Cotton gravity is constructed from the Cotton tensor, and it has been shown that one can rewrite the field equations of Cotton gravity as a Codazzi equation $\nabla_{[a}C_{b]c}=0$ [34; 53]; in this regard, Cotton gravity is in the same class as the earlier theory of Cook and Chen, as we will show that the field equations for the latter can also be reformulated as a Codazzi equation. Conformal Killing gravity, has a different structure altogether, as the field equations resemble the defining equation for a conformal Killing tensor.
@@
Cotton gravity was recently the focus of a debate [15; 51; 16; 52] over its viability as a physical theory, as it was discovered that the symmetry-reduced equations of Cotton gravity are underdetermined. It is been argued in [51; 52] that while the equations of Cotton gravity are indeed underdetermined in cases of high symmetry (for instance, cosmological or spherically symmetric spacetimes), this is not necessarily the case in generic spacetimes. Similar concerns apply to the theory of Cook and Chen, as they share the same Codazzi structure. Another concern that applies to Codazzi type theories and conformal Killing gravity, is the fact that they are only defined at the level of the field equations; these theories suffer from the lack of a complete variational principle, as the variational principles proposed so far require holding some of the variables (such as the spacetime metric $g_{ab}$) fixed. One difficulty is that the energy-momentum tensor is obtained from variations of the metric tensor in the matter sector, but derivatives of the energy-momentum tensor appear in the field equations.
As remarked earlier, the primary aim of this article is to assess the viability of the aforementioned approaches to decoupling the CC from the vacuum energy (without introducing additional constraints). We focus on two specific requirements. First, we require that the field equations can be reformulated in a manner that yields a well-posed initial value problem—in doing so, we aim to shed light on the recent debate over the Codazzi equation. Second, we propose variational principles (including matter) for the trace-free Einstein equations, the Codazzi equation, and the conformal Killing equation without introducing additional constraints. It should be noted that some variational principles for the trace-free Einstein equations were recently proposed [36; 37] that avoid additional constraints, but our approach is the only one so far to explicitly incorporate the matter sector while avoiding the need for fine-tuning. Though our proposed actions are not intended to be fundamental, as they require the introduction of auxiliary fields and ghosts (which do not exhibit classical pathologies), they may serve as a first step toward a more complete theory.
2 Theories
The theories we consider here differ significantly in their construction, but they all share the property that they yield the unconstrained Einstein equations in a manner that is entirely independent of the vacuum energy contribution in the energy-momentum tensor (avoiding the need for fine-tuning), with the cosmological constant term arising as a constant of integration. In this section, we briefly review the aforementioned theories.
2.1 Trace-free Einstein equations
@@
The trace-free Einstein equations may be obtained by assuming that the determinant of the metric is held fixed under the variation of the action; in this form, the theory is called unimodular gravity. However, the equations themselves can be written without assuming such a constraint [24; 13]:
@@
2.2 Bianchi gravity
The works [17; 14] consider a field theory in which the connection and curvature tensors are treated as analogues of the respective gauge field and field strength tensor of electrodynamics, yielding:
@@ 2.2 Bianchi gravity
Equation (9) has the form of a Codazzi equation [34; 53], which is of recent interest in the Cotton gravity program [26].
2.3 Cotton gravity
Cotton gravity is based on the Cotton tensor:
@@ 2.3 Cotton gravity
one can write the field equation for Cotton gravity as:
@@
2.4 ‘Conformal” Killing gravity
The equations for the so-called conformal Killing gravity theory proposed in [27] may be written in the form [35]:
@@
where $u_{a}$ is a one-form; $K_{ab}$ is a symmetric Killing tensor when $u_{a}=0$. However, as pointed out in [6; 5] a simple redefinition $\bar{K}_{ab}:=6K_{ab}-g_{ab}K$ reveals that any nontrivial solution of Eq. (16) requires the existence of a Killing tensor given by $\bar{K}_{ab}$. It follows that the solutions of Eq. (16) either trivialize to $K_{ab}=-\lambda g_{ab}$, or admit a Killing tensor. For this reason, it is perhaps more appropriate to refer to this theory as “Killing tensor gravity”, but we retain here the common terminology in the literature. Moreover, we note that perturbations of Killing tensor solutions satisfying $K_{ab}\neq-\lambda g_{ab}$ must also admit a Killing tensor; since our universe (being inhomogeneous and anisotropic on small scales) does not have the required symmetries to admit a global Killing tensor, we regard such solutions as unphysical.
3 Codazzi equation
Equations (9) and (14) have the form of a Codazzi equation:
@@ 3 Codazzi equation
where $C_{ab}=C_{ba}$ is termed a Codazzi tensor, following the terminology of [34; 53]. One can perhaps imagine a class of theories defined by an equation of the form in Eq. (18). Now there is some debate over whether equations of this type are predictive [15; 51; 16; 52]; in particular, it has been claimed in [16] that the Codazzi equations are underdetermined, and in a responding comment it was claimed that this only occurs for highly symmetric spacetimes [52]. In the following, we attempt to shed some light on this dispute.
@@
Of course, Eq. (20) is underdetermined due to the term $\nabla_{a}{\nabla}_{c}C{}^{cb}$. In an orthonormal basis, the derivative $\partial^{2}_{t}C{{}_{t}}^{b}$ disappears from Eq. (20), so that the time evolution of the components $C{{}_{t}}^{b}$ is ill-defined—the equations are indeed underdetermined. However, Eq. (20) reveals a straightforward resolution; one should supply the Codazzi equation with a constraint equation, and a natural one is the divergence-free condition $\nabla_{a}C^{ab}=0$, which is essentially a statement that $C^{ab}$ satisfies a local conservation law. [^2] This is natural in the sense that each term in Eq. (15) for $C_{ab}$ independently satisfies a conservation law when the Einstein equation holds (up to a cosmological term, also cf. Eq. (8)) and in particular when the energy-momentum tensor satisfies the local conservation law $\nabla_{a}T^{ab}=0$. Of course, an attempt to directly solve the Codazzi equations supplemented with the constraint $\nabla_{a}C^{ab}=0$ may still lead to ambiguities, but such ambiguities correspond to a partial freedom (which we restrict in the next paragraph) in choosing initial data for $C_{ab}$ and $\partial_{t}C_{ab}$ for the system:
@@
While solutions of Eq. (18) satisfying the divergence-free constraint will also satisfy the system in Eq. (22), there is no guarantee that solutions of Eq. (22) will satisfy Eq. (18). To identify the conditions on the initial data that allow one to recover solutions of the Codazzi equation, we consider the application of the D’Alembertian to the following quantity:
@@
Note that the right-hand side contains at most first derivatives of $C_{ab}$ (and also that the divergence-free condition removes the first term). One can then regard Eq. (24) as a wave equation for $Q_{cde}$, with $C_{ab}$ being an independent variable satisfying Eq. (20) and the constraint (24). Now consider initial data consistent with Eq. (18), so that $Q_{cde}=0$ (but not its derivatives). For nonvanishing curvature, the right-hand side will generically produce a source unless $C_{ab}=\lambda g_{ab}$ on a Cauchy hypersurface and $\nabla_{c}C_{ab}=0$. This indicates that if $C_{ab}\neq\lambda g_{ab}$ and $\nabla_{c}C_{ab}\neq 0$ satisfies Eq. (24) on a Cauchy surface, the time evolution of $C_{ab}$ under Eq. (20) will in general generate violations of Eq. (24). There may also be special cases (highly symmetric spacetimes, for instance) in which Eq. (18) may admit nontrivial solutions $C_{ab}\neq\lambda g_{ab}$, but one cannot expect the existence of solutions to Eq. (18) for generic initial data. Observe that Eq. (24) and the conclusions outlined in the previous paragraph are independent of the constraint $\nabla_{a}C^{ab}=0$.
To summarize, the Codazzi equation, supplied with the divergence-free constraint $\nabla_{a}C^{ab}=0$, implies that $C_{ab}$ satisfies Eq. (22), which has a well-defined initial value problem. While Eq. (22) also admits solutions that do not satisfy the Codazzi equation, one can restrict the space of initial data to eliminate such solutions that remain. The preceding analysis surrounding (24) indicates that in a general spacetime, one can only guarantee the existence of solutions to the Codazzi equation for initial data also satisfying $C_{ab}=\lambda g_{ab}$ and $\nabla_{c}C_{ab}=0$, where $\lambda$ is an arbitrary constant; other solutions will likely require background geometries with special properties, as argued in [51; 52]. We find that both sets of authors, [15; 16] and [51; 52], present valid points: theories based on the Codazzi tensors are indeed underdetermined, but the space of solutions is nonetheless highly restricted in generic settings.
4 Variational principle
The variational principles presented in the literature for the trace-free Einstein equations [58] and theories admitting a Codazzi equation formulation (Bianchi gravity [17; 14] and Cotton gravity [26]) typically require holding certain quantities fixed; for the trace-free Einstein equations, the metric determinant is held fixed (corresponding to an additional constraint), and for the Codazzi equation, the metric is held fixed (as variations in the metric will generally introduce additional unwanted constraints). We therefore seek a general variational principle that yields the desired field equations (including contributions from the matter sector) without unwanted constraints under a full variation—that is, a variation with respect to all variables appearing in the action.
@@ 4 Variational principle
with $\underline{R}_{ab}:=R_{ab}-\tfrac{1}{4}Rg_{ab}$ and $\underline{Z}_{ab}:=R_{ab}-\tfrac{1}{4}Zg_{ab}$. We emphasize that no constraint on the metric determinant is required.
@@
where $\phi$ is a matter field (assumed to be nonspinorial and minimally coupled), $\chi$ is an extra field that is equal to $\phi$ on shell; we note that $\chi$ is necessarily a ghost, but we argue later that classically, it does not exhibit runaway behavior. Now $h^{ab}$ and $J$ are auxiliary fields, $\zeta$ is a Lagrange multiplier [^4] In particular, $\zeta$ is defined so that $J\cdot\zeta\cdot J$ is a scalar., and the following have been defined:
@@
5 Discussion
In this article, we considered some theories in which the CC arises as an integration constant and yields the Einstein equation without additional constraints. This class includes trace-free Einstein gravity, the Codazzi equation, and conformal Killing gravity. An interesting question, which we have not considered, is whether there are other theories that satisfy this property; we have made an attempt to consider all the proposals in the literature, but the existence and construction of other theories sharing this property (which may perhaps avoid the properties discussed in the following paragraphs) is an interesting question for future investigation.